OpenStudy (anonymous):

Pretty sure I am over thinking this question! Help please :) Given the function f(x) = x + 3x + 4 , find the value of f(–1). f(–1) = 8 f(–1) = –2 f(–1) = 0 f(–1) = 2

OpenStudy (anonymous):

\[f(-1)=-1-3+4=0\]...assuming you didn't mean \[f(x)=x^2+3x+4\]

OpenStudy (ddcamp):

Just replace each x with (-1): f(-1) = (-1) + 3(-1) + 4 f(-1) = -1 - 3 + 4 f(-1) = 0

OpenStudy (anonymous):

i assume this is \[f(x)=x^2+3x+4\] think of it as \[f(\diamondsuit)=\diamondsuit ^2+3\times \diamondsuit +4\]

OpenStudy (anonymous):

replace \(\diamondsuit\) by \(-1)\) with judicious use of parentheses and compute \[f(-1)=(-1)^2+3\times (-1)+4\]

OpenStudy (anonymous):

thanks a bunch guys :)